Two professors at a local college developed a new teaching curriculum designed to increase students grades in math classes.(a) State the appropriate null and alternative hypotheses. 0.56 Ho: P = 0.56 versus Hy: p > (Type integers or decimals. Do noSOLVE (D) AND SHOW HOW YOU GOT THE ANSWER PLEASE!
Answer #1

d)

here this is binomial with parameter n=15 and p=0.56

p value =P(getting 11 or more with grade A, B and C if p value =0.56)

=P(X=11)+P(X=12)+P(X=13)+P(X=14)+P(X=15)

= (15C11)*(0.56)^11*(0.44)^4+(15C12)*(0.56)^12*(0.44)^3+(15C13)*(0.56)^13*(0.44)^2+(15C14)*(0.56)^14*(0.44)^1+(15C15)*(0.56)^15*(0.44)^0

=0.0869+0.0369+0.0108+0.0020+0.0002

=0.137